# Differentially Private Clustering in Data Streams

### v2

@@ 5 DP Merge And Reduce Algorithm
1:Point $x^{\prime}_{t}$ to be added to coreset, DP parameters $\varepsilon,\varepsilon_{1},\delta_{1}$
2:Initialize($\varepsilon,\varepsilon_{1},\delta_{1}$):
3:Block size $M$$\triangleright$ Value of $M$ depends on our choice of DP semicoreset Algorithm. See Corollary 14 and Corollary 16 for exact values
4:$P_{0}=\emptyset$
5:Size of $P_{0}$ as ${p}_{0}=0$
6$\triangleright$ We only store $P_{0}$ in the actual algorithm, the rest of the $P_{i}$’s are only used in the accuracy analysis, so we do not run the following line in the actual algorithm. $N_{r}$ denotes the number of points in ring $R_{r}$.
7:$P_{i}=\emptyset$ for all $i=1,\ldots,u$ where $u=\lfloor\log(2{N}_{r}/M)\rfloor+1$
8:$\gamma_{i}:=\gamma/Ci^{2}$ for $i=1,\ldots,u-1$
9:Update($x^{\prime}_{t}$):
10:if $x^{\prime}_{t}\neq\perp$ then
11:     Insert $x^{\prime}_{t}$ into $P_{0}$
12:     $p_{0}=p_{0}+1$$\triangleright$ update the size of $P_{0}$
[…]
14:         $Q_{1}\leftarrow{\mathcal{A}}(P_{0},k,d,\varepsilon,\delta,\kappa,\eta_{1},\eta_{2})$triangleright ${\mathcal{A}}$ computes a $(\varepsilon_{1},\delta_{1})$-DP $(\kappa,\eta_{1},\eta_{2})$-semicoreset, e.g., [42],[29]
15:         $P_{1}=P_{1}\cup P_{0}$
16:         Delete elements in $P_{0}$, and set ${p}_{0}=0$
17:         $i=1$
18:         while $\exists\ Q^{\prime}_{i}$ such that $Q_{i}$ and $Q^{\prime}_{i}$ are both at level $i$ do
19:              $\triangleright$ Merge-Reduce computes $(1+\gamma_{i})$-coreset $Q_{i+1}$ of $Q_{i}\cup Q^{\prime}_{i}$ and deletes $Q_{i},Q^{\prime}_{i}$
20:              $Q_{i+1}=\textbf{Merge\text{-}Reduce}({\mathcal{B}},Q_{i},Q^{\prime}_{i},\gamma_{i})$triangleright ${\mathcal{B}}$ computes a non-DP $(1+\gamma_{i})$-coreset, e.g., [19]
21:              $P_{i+1}=P_{i+1}\cup P_{i}$triangleright Lines 21 and 22 are for analysis purpose only
22:              Delete elements in $P_{i}$
23:              $i=i+1$
24:         return $\textbf{Merge\text{-}Reduce}({\mathcal{B}},\{Q_{i}\}_{i\leq u},\gamma/3)$
25:     else
26:         return $\emptyset$
27:else
28:     return $\emptyset$
@@
1:Size of set $P_{0}$ denoted as $p_{0}$, Block size $M$, Privacy parameter $\varepsilon$
2$\top$ if the size of set $P_{0}$ exceeds noisy threshold and $\perp$ otherwise
3:$\hat{M}=M+{\textsf{Lap}}(2/\varepsilon)$
4:$\nu={\textsf{Lap}}(4/\varepsilon)$
5:if $p_{0}+{\nu}\geq\hat{M}$ then
6:     return $\top$
7:else
8:     return $\bot$
@@
$\geq(1-\gamma)\textsf{cost}(C,P)-\eta_{1}\textsf{cost}(C^{opt}_{P},P)-\eta_{2}+(1-\gamma)\textsf{cost}(C,P^{\prime})-\eta_{1}\textsf{cost}(C^{opt}_{P^{\prime}},P^{\prime})-\eta_{2}$
@@ 5.1 Proof of Theorem 10
With probability at least $1-\xi-\xi_{A}-\xi_{B}$, DP-Merge-Reduce framework releases a $((1+\gamma)\kappa,(1-\gamma)\eta_{1},(\frac{4N}{M}-1)(1+\gamma)\eta_{2}+\tilde{M})$-semicoreset of $P$. Where $\tilde{M}:={M}+\frac{6}{\varepsilon}\log(\frac{2T}{\xi})$.
@@
$\mathsf{cost}(C,Q_{i}\cup Q^{\prime}_{i})$ $\geq\kappa\prod^{i-1}_{j=1}(1-\gamma_{j})\mathsf{cost}(C,P_{i}\cup P^{\prime}_{i})-\prod^{i-1}_{j=1}(1-\gamma_{j})\eta_{1}\mathsf{cost}(C^{opt}_{P_{i}\cup P^{\prime}_{i}},P_{i}\cup P^{\prime}_{i})-\prod^{i-1}_{j=1}(1-\gamma_{j})2^{i}\eta_{2}$ (35)
@@
Note that if ${p}_{0}+\nu<\hat{M}$ then we do not release anything. So we have to account for an additional additive error of ${M}+\frac{6}{\varepsilon}\log(\frac{2T}{\xi})$ w.p. $1-\xi$ in this case. If ${p}_{0}+\nu\geq\hat{M}$, then we proceed by computing a DP coreset using Algorithm ${\mathcal{A}}$ with failure probability $\xi_{A}$. We also compute all the coresets past the first level using Algorithm ${\mathcal{B}}$ and a failure probability of $\xi_{B}/2u$, where $u$ is the number of levels. Thus for a fixed run of DP-Merge-Reduce, by a union bound, the total failure probability for this part is at most $\xi_{B}$. ∎
@@
Claim 9.
If $p_{0}\geq\frac{3M}{2}$ then with probability $1-\xi$, ${p}_{0}+\nu\geq\hat{M}$.
@@
Thus, we only need to store at most $23M/$ 2 points for the block of input points at level 0, plus the additional space required to execute the coreset construction. Note that the semicoreset computation of level 0 blocks consumes space $S_{\mathcal{A}}(M,k,d,\varepsilon,\delta,\kappa,\eta_{1},\eta_{2})$ and since the largest coreset construction wrt non-DP algorithm ${\mathcal{B}}$ is the union of at most $u$ coresets that is reduced to a single coreset, and the largest input to non-DP coreset algorithm is the resulting DP semicoreset size — the additional storage size is at most $S_{\mathcal{B}}(SZ_{\mathcal{A}}(M,k,d,\varepsilon,\delta,\kappa,\eta_{1},\eta_{2}),k,d,\gamma)+(u-1)\cdot S_{\mathcal{B}}(SZ_{\mathcal{B}}(M,k,d,\gamma),k,d,\gamma)$. Note that the resulting coreset size for non-DP coreset algorithm ${\mathcal{B}}$ is independent of the input set size.
@@
Appendix A Proofs for Bicriteria Approximation
Proof of Claim 2.

### v3

@@ Appendix C DP Merge And Reduce Algorithm
Data: Input point $x^{\prime}_{t}$ to be added to semicoreset, DP parameters $\varepsilon,\varepsilon_{1},\delta_{1}$
Initialize($\varepsilon,\varepsilon_{1},\delta_{1}$):
Block size $M$;
$\triangleright$ Value of $M$ depends on our choice of DP clustering algorithm. See Theorem 2 and Theorem 3 for exact values
$P_{0}=\emptyset$;
Size of $P_{0}$ as ${p}_{0}=0$;
$\triangleright$ We only store $P_{0}$ in the actual algorithm, the rest of the $P_{i}$’s are only used in the accuracy analysis, so we do not run the following line in the actual algorithm. $N_{r}$ denotes the number of points in ring $R_{r}$.
$P_{i}=\emptyset$ for all $i=1,\ldots,u$ where $u=\lfloor\log(2{N}_{r}/M)\rfloor+1$;
$\gamma_{i}:=\gamma/Ci^{2}$ for $i=1,\ldots,u-1$;
0.75em Update($x^{\prime}_{t}$):
if $x^{\prime}_{t}\neq\perp$ then
  Insert $x^{\prime}_{t}$ into $P_{0}$;
 $p_{0}=p_{0}+1$ $\triangleright$ update the size of $P_{0}$;
 if $\textbf{LevelZero-AboveThreshold}(p_{0},M,\varepsilon)\to\top$ then
    $\triangleright$ See algorithm 5
    $Q_{1}\leftarrow{\mathcal{A}}(P_{0},k,d,\varepsilon,\delta,\kappa,\eta_{1},\eta_{2})$ $\triangleright$ $\mathcal{A}$} computes a $(\varepsilon_{1},\delta_{1})$-DP $(\kappa,\eta_{1},\eta_{2})$-semicoreset;
    $P_{1}=P_{1}\cup P_{0}$;
     Delete elements in $P_{0}$, and set ${p}_{0}=0$;
    $i=1$;
    while $\exists\ Q^{\prime}_{i}$ such that $Q_{i}$ and $Q^{\prime}_{i}$ are both at level $i$ do
       $\triangleright$ Merge-Reduce computes $(1+\gamma_{i})$-coreset $Q_{i+1}$ of $Q_{i}\cup Q^{\prime}_{i}$ and deletes $Q_{i},Q^{\prime}_{i}$
       $Q_{i+1}=\textbf{Merge\text{-}Reduce}({\mathcal{B}},Q_{i},Q^{\prime}_{i},\gamma_{i})$;
       $\triangleright$ $\mathcal{B}$} computes a non-DP $(1+\gamma_{i})$-coreset, e.g., [20]
       $P_{i+1}=P_{i+1}\cup P_{i}$;
       $\triangleright$ Current line and the one above are for analysis purpose only
        Delete elements in $P_{i}$;
       $i=i+1$;
     return $\textbf{Merge\text{-}Reduce}({\mathcal{B}},\{Q_{i}\}_{i\leq u},\gamma/3)$ ;
 else
      return $\emptyset$;
else
   return $\emptyset$;
@@
Data: Size of set $P_{0}$ denoted as $p_{0}$, Block size $M$, Privacy parameter $\varepsilon$
Result: $\top$ if the size of set $P_{0}$ exceeds noisy threshold and $\perp$ otherwise
$\hat{M}=M+{\textsf{Lap}}(2/\varepsilon)$;
$\nu={\textsf{Lap}}(4/\varepsilon)$;
if $p_{0}+{\nu}\geq\hat{M}$ then
   return $\top$;
else
   return $\bot$;
@@
$\geq(1-\gamma)\textsf{cost}(C,P)-\eta_{1}\textsf{cost}(C^{opt}_{P},P)-\eta_{2}$
$+(1-\gamma)\textsf{cost}(C,P^{\prime})-\eta_{1}\textsf{cost}(C^{opt}_{P^{\prime}},P^{\prime})-\eta_{2}$
@@ C.1 Proof of Theorem 5
[…]
@@
$\mathsf{cost}(C,Q_{i}\cup Q^{\prime}_{i})$ $\geq\kappa\prod^{i-1}_{j=1}(1-\gamma_{j})\mathsf{cost}(C,P_{i}\cup P^{\prime}_{i})$
$-\prod^{i-1}_{j=1}(1-\gamma_{j})\eta_{1}\mathsf{cost}(C^{opt}_{P_{i}\cup P^{\prime}_{i}},P_{i}\cup P^{\prime}_{i})-\prod^{i-1}_{j=1}(1-\gamma_{j})2^{i}\eta_{2}$ (8)
@@
Note that if $^{p}_{0}+\nu<\hat{M}$ then we do not release anything. So we have to account for an additional additive error of ${M}+\frac{6}{\varepsilon}\log(\frac{2T}{\xi})$ w.p. $1-\xi$ in this case. If $^{p}_{0}+\nu\geq\hat{M}$, then we proceed by computing a DP coreset using Algorithm $\mathcal{A}$} with failure probability $\xi_{A}$. We also compute all the coresets past the first level using Algorithm $\mathcal{B}$} and a failure probability of $\xi_{B}/2u$, where $u$ is the number of levels. Thus for a fixed run of DP-Merge-Reduce, by a union bound, the total failure probability for this part is at most $\xi_{B}$. ∎
@@
Lemma 17.
If $p_{0}\geq\frac{3M}{2}$ then with probability $1-\xi$, $^{p}_{0}+\nu\geq\hat{M}$.
@@
Thus, we only need to store at most $3M/2$ points for the block of input points at level 0, plus the additional space required to execute the coreset construction. Note that the semicoreset computation of level 0 blocks consumes space $S_{\mathcal{A}}(M,k,d,\varepsilon,\delta,\kappa,\eta_{1},\eta_{2})$ and since the largest coreset construction wrt non-DP algorithm $\mathcal{B}$} is the union of at most $u$ coresets that is reduced to a single coreset, and the largest input to non-DP coreset algorithm is the resulting DP semicoreset size — the additional storage size is at most $S_{\mathcal{B}}(SZ_{\mathcal{A}}(M,k,d,\varepsilon,\delta,\kappa,\eta_{1},\eta_{2}),k,d,\gamma)+(u-1)\cdot S_{\mathcal{B}}(SZ_{\mathcal{B}}(M,k,d,\gamma),k,d,\gamma)$. Note that the resulting coreset size for non-DP coreset algorithm $\mathcal{B}$} is independent of the input set size.
@@
Appendix D Missing Proofs from Section 2
D.1 Space and Size
Lemma 19.
algorithm 1 consumes
$+\tilde{O}(k\log(T/M)\log(\Lambda)\log k\cdot\gamma^{-4})+O({k}\log(k)\log^{2}(\Lambda)\log T)$
$+O(k\log(\Lambda)\log^{2}(k)\operatorname*{poly}\left(\log\left({T}{\Lambda}k\right)\right))$
space.
Proof.
We focus on the last term in eq. 13. In particular, since the space used by $\mathcal{B}$} is linear in the coreset size, we have that $S_{\mathcal{B}}(SZ_{\mathcal{B}}(\cdot),k,d,\gamma)=\tilde{O}(k\log k\cdot\gamma^{-4})$. Next we simplify the adjoining sum
Lemma 20.
Proof.
D.2 Accuracy
We analyze the accuracy of algorithm 1 in the sequel. We first state the accuracy guarantee of the DP semicoreset $\hat{{\mathcal{Y}}}^{(t)}_{r}$ released by $\textsf{DP-Merge-Reduce}_{r}$ for each ring $R^{(t)}_{r}$ at timestep $t\in[T]$ in Lemma 21. We defer the detailed proofs to Appendix D.
Lemma 21.
[…]
$\leq\mathsf{cost}({\mathcal{C}},\hat{{\mathcal{Y}}}^{(t)}_{r})$
Proof.
Theorem 13.
$\frac{(1-\gamma)^{2}}{\kappa}\mathsf{cost}({\mathcal{C}},{\mathcal{S}})-((1-\gamma)^{2}\eta_{1}\Lambda^{2}+16C_{M}(1-\gamma^{2}))\cdot\mathsf{cost}({\mathcal{C}}^{opt}_{\mathcal{S}},{\mathcal{S}})$ (16)
$-C_{M}(1-\gamma^{2})\cdot V^{\prime}(d,k,\varepsilon,T,\Lambda))-(1-\gamma)V^{\prime\prime}(d,k,\varepsilon,T,\Lambda)$
[…]
$\leq(1+\gamma)^{2}\kappa\mathsf{cost}({\mathcal{C}},{\mathcal{S}})+16C_{M}(1+\gamma)^{2}\mathsf{cost}({\mathcal{C}}^{opt}_{\mathcal{S}},{\mathcal{S}})$
$+C_{M}(1+\gamma)^{2}\cdot V^{\prime}(d,k,\varepsilon,T,\Lambda))+(1+\gamma)V^{\prime\prime}(d,k,\varepsilon,T,\Lambda)$ (17)
Proof.
Fix an epoch $T_{i}$. Let $\hat{{\mathcal{Y}}}^{(t)}$ be the output of algorithm 1 at timestep $t\in T_{i}$. Since the centers in ${\mathcal{F}}|_{T_{i}}$ and consequently the rings $R_{r}$ are fixed, we can compute the cost of $\hat{{\mathcal{Y}}}^{(t)}$ by summing Lemma 21 from Lemma 21 over all rings $R_{r}$ where $1\leq r\leq\log(\Lambda)$ as follows:
$\frac{(1-\gamma)}{\kappa}\sum_{r}\mathsf{cost}({\mathcal{C}},R^{(t)}_{r})-(1-\gamma)\eta_{1}\sum_{r}\textsf{cost}({\mathcal{C}}^{opt}_{R^{(t)}_{r}},R^{(t)}_{r})\cdot(2^{r})^{2}$
$-(1+\gamma)\eta_{2}\sum_{r}(\frac{4N^{(t)}_{r}}{M}-1)\cdot(2^{r})^{2}-\sum_{r}\tilde{M}\cdot(2^{r})^{2}$
$\leq\mathsf{cost}({\mathcal{C}},\hat{{\mathcal{Y}}}^{(t)})$
Next, we take the union of $\hat{{\mathcal{Y}}}^{(t)}$ over all timesteps in epoch $T_{i}$. Note that for a fixed epoch and fixed ring $R_{r}$, the additive error $\tilde{M}$ is incurred at most once since this error stems from privately testing if the number of points at the base level is larger than the block size $M$ (see algorithm 5) in $\textsf{DP-Merge-Reduce}_{r}$. Thus summing Section D.2 over all timesteps in $T_{i}$ gives us
$\frac{(1-\gamma)}{\kappa}\mathsf{cost}({\mathcal{C}},{\mathcal{S}}|_{T_{i}})-(1-\gamma)\eta_{1}\Lambda^{2}\textsf{cost}({\mathcal{C}}^{opt}_{{\mathcal{S}}|_{T_{i}}},{\mathcal{S}}|_{T_{i}})$
$-(1+\gamma)\eta_{2}\sum_{t\in T_{i}}\sum_{r}(\frac{4N^{(t)}_{r}}{M}-1)\cdot(2^{r})^{2}-\tilde{M}\cdot\frac{4}{3}(\Lambda^{2}-1)$
$\leq\mathsf{cost}({\mathcal{C}},\hat{{\mathcal{Y}}}^{(t)}|_{t\in T_{i}})$
Lemma 22.
If $x_{i}$ is assigned to the ring $R_{r}$, let $r(x_{i}):=2^{r}$. Then
Proof.
Observe that $\sum_{t\in T_{i}}\sum_{r}N^{(t)}_{r}\cdot(2^{r})^{2}=\sum_{x\in{\mathcal{S}}|_{T_{i}}}r(x)^{2}$, thus we can use Lemma 22 to simplify the additive error in Section D.2 to obtain
$\frac{(1-\gamma)}{\kappa}\mathsf{cost}({\mathcal{C}},{\mathcal{S}}|_{T_{i}})-(1-\gamma)\eta_{1}\Lambda^{2}\textsf{cost}({\mathcal{C}}^{opt}_{{\mathcal{S}}|_{T_{i}}},{\mathcal{S}}|_{T_{i}})$
$-\frac{4}{M}(1+\gamma)\eta_{2}(4\alpha\cdot\mathsf{cost}({\mathcal{C}}^{opt}_{{\mathcal{S}}|_{T_{i}}},{\mathcal{S}}|_{T_{i}})+V(d,k,\varepsilon,T,\Lambda))-\tilde{M}\cdot\frac{4\Lambda^{2}}{3}$
$\leq\mathsf{cost}({\mathcal{C}},\hat{{\mathcal{Y}}}^{(t)}|_{t\in T_{i}})$
$\frac{(1-\gamma)^{2}}{\kappa}\mathsf{cost}({\mathcal{C}},{\mathcal{S}}|_{T_{i}})-(1-\gamma)^{2}\eta_{1}\Lambda^{2}\textsf{cost}({\mathcal{C}}^{opt}_{{\mathcal{S}}|_{T_{i}}},{\mathcal{S}}|_{T_{i}})$
$-\frac{4}{M}(1-\gamma^{2})\eta_{2}(4\alpha\cdot\mathsf{cost}({\mathcal{C}}^{opt}_{{\mathcal{S}}|_{T_{i}}},{\mathcal{S}}|_{T_{i}})+V(d,k,\varepsilon,T,\Lambda))-\frac{4\Lambda^{2}(1-\gamma)\tilde{M}}{3}$
$\leq\mathsf{cost}({\mathcal{C}},\hat{{\mathcal{Y}}}_{T_{i}})$
$\leq(1+\gamma)^{2}\kappa\mathsf{cost}({\mathcal{C}},{\mathcal{S}}|_{T_{i}})+\frac{4}{M}(1+\gamma)^{2}\eta_{2}(4\alpha\cdot\mathsf{cost}({\mathcal{C}}^{opt}_{{\mathcal{S}}|_{T_{i}}},{\mathcal{S}}|_{T_{i}})$
$+V(d,k,\varepsilon,T,\Lambda))+\frac{4\Lambda^{2}(1+\gamma)\tilde{M}}{3}$ (22)
Observe that the total number of epochs is bounded by the total number of centers in $\mathcal{F}$} at the end of the stream. By Theorem 4 we know that $|{\mathcal{F}}|=O({k}\log^{2}(k)\log(\Lambda)\log T)$. Therefore, we can sum Section D.2 over all epochs to have that for any set of $k$ centers $\mathcal{C}$}:
$\frac{(1-\gamma)^{2}}{\kappa}\mathsf{cost}({\mathcal{C}},{\mathcal{S}})-(1-\gamma)^{2}\eta_{1}\Lambda^{2}\textsf{cost}({\mathcal{C}}^{opt}_{{\mathcal{S}}},{\mathcal{S}})-\frac{4}{M}(1-\gamma^{2})\eta_{2}(4\alpha\cdot\mathsf{cost}({\mathcal{C}}^{opt}_{\mathcal{S}},{\mathcal{S}})$
$+V^{*}(d,k,\varepsilon,T,\Lambda))-{|{\mathcal{F}}|}\cdot\frac{4\Lambda^{2}(1-\gamma)\tilde{M}}{3}$
[…]
$\leq(1+\gamma)^{2}\kappa\mathsf{cost}({\mathcal{C}},{\mathcal{S}})+\frac{4}{M}(1+\gamma)^{2}\eta_{2}(4\alpha\cdot\mathsf{cost}({\mathcal{C}}^{opt}_{\mathcal{S}},{\mathcal{S}})$
$+V^{*}(d,k,\varepsilon,T,\Lambda))+{|{\mathcal{F}}|}\cdot\frac{4\Lambda^{2}(1+\gamma)\tilde{M}}{3}$ (23)
Grouping like terms together, we have
$\frac{(1-\gamma)^{2}}{\kappa}\mathsf{cost}({\mathcal{C}},{\mathcal{S}})-((1-\gamma)^{2}\eta_{1}\Lambda^{2}+\frac{16\alpha}{M}(1-\gamma^{2})\eta_{2})\cdot\mathsf{cost}({\mathcal{C}}^{opt}_{\mathcal{S}},{\mathcal{S}})$
$-\frac{4}{M}(1-\gamma^{2})\eta_{2}\cdot V^{*}(d,k,\varepsilon,T,\Lambda))-{|{\mathcal{F}}|}\cdot\frac{4\Lambda^{2}(1-\gamma)\tilde{M}}{3}$
[…]
$\leq(1+\gamma)^{2}\kappa\mathsf{cost}({\mathcal{C}},{\mathcal{S}})+\frac{4}{M}(1+\gamma)^{2}\eta_{2}(4\alpha\cdot\mathsf{cost}({\mathcal{C}}^{opt}_{\mathcal{S}},{\mathcal{S}})$
$+V^{*}(d,k,\varepsilon,T,\Lambda))+{|{\mathcal{F}}|}\cdot\frac{4\Lambda^{2}(1+\gamma)\tilde{M}}{3}$ (24)
We set
$M:=\frac{\alpha\eta_{2}}{C_{M}}$ (25)
where $C_{M}$ is a parameter chosen in the sequel (see remark Remark). Simplifying Section D.2 and taking a union bound over all rings and epochs, the desired claim in the theorem statement holds with probability $1-\log(\Lambda)|{\mathcal{F}}|3\xi-\frac{1}{k^{O(\operatorname*{poly}(k,\log(\Lambda))}}$. Thus we set $\xi:=\frac{1}{3|{\mathcal{F}}|\log(\Lambda)T^{2}}$. ∎
Theorem 14 (Main Accuracy).
Given dimension $d$, clustering parameter $k$, arbitrary parameter $C_{M}$, non-DP $(1+\gamma)$-coreset, $O(d^{3})$-approximate bicriteria solution from Theorem 4, DP $(\kappa,\eta_{1},\eta_{2})$-semicoreset, and privacy parameter $\varepsilon$. Let $^{\mathcal{C}}_{\hat{{\mathcal{Y}}}}$ be the set of $k$ centers obtained from running the offline $\rho$-approx non-DP $k$-means algorithm on $\hat{{\mathcal{Y}}}$. Then,
$\mathsf{cost}({\mathcal{C}}_{\hat{{\mathcal{Y}}}},{\mathcal{S}})$
Proof.
$\frac{(1-\gamma)^{3}}{\kappa}\mathsf{cost}({\mathcal{C}},{\mathcal{S}})-((1-\gamma)^{3}\eta_{1}\Lambda^{2}+16C_{M}(1-\gamma^{2})(1-\gamma))\cdot\mathsf{cost}({\mathcal{C}}^{opt}_{\mathcal{S}},{\mathcal{S}})$
$-C_{M}(1-\gamma^{2})(1-\gamma)\cdot V^{\prime}(d,k,\varepsilon,T,\Lambda))-(1-\gamma)^{2}V^{\prime\prime}(d,k,\varepsilon,T,\Lambda)$
$\leq\mathsf{cost}({\mathcal{C}},\hat{{\mathcal{Y}}}_{off})$
$\leq(1+\gamma)^{3}\kappa\mathsf{cost}({\mathcal{C}},{\mathcal{S}})+16C_{M}(1+\gamma)^{3}\mathsf{cost}({\mathcal{C}}^{opt}_{\mathcal{S}},{\mathcal{S}})$
$+C_{M}(1+\gamma)^{3}\cdot V^{\prime}(d,k,\varepsilon,T,\Lambda))+(1+\gamma)^{2}V^{\prime\prime}(d,k,\varepsilon,T,\Lambda)$ (27)
Recall from the theorem statement that $^{\mathcal{C}}_{\hat{{\mathcal{Y}}}}$ is the set of $k$ centers obtained from running the offline $\rho$-approx non-DP $k$-means algorithm on $\hat{{\mathcal{Y}}}_{off}$ which gives us the following guarantee,
Since Section D.2 is true for any set of $k$ centers, in particular it holds for $^{\mathcal{C}}_{\hat{{\mathcal{Y}}}}$, which gives us the following
$\frac{(1-\gamma)^{3}}{\kappa}\mathsf{cost}({\mathcal{C}}_{\hat{{\mathcal{Y}}}},{\mathcal{S}})-((1-\gamma)^{3}\eta_{1}\Lambda^{2}+16C_{M}(1-\gamma^{2})(1-\gamma))\cdot\mathsf{cost}({\mathcal{C}}^{opt}_{\mathcal{S}},{\mathcal{S}})$
$-C_{M}(1-\gamma^{2})(1-\gamma)\cdot V^{\prime}(d,k,\varepsilon,T,\Lambda))-(1-\gamma)^{2}V^{\prime\prime}(d,k,\varepsilon,T,\Lambda)$
$\leq\mathsf{cost}({\mathcal{C}}_{\hat{{\mathcal{Y}}}},\hat{{\mathcal{Y}}}_{off})$
$\leq(1+\gamma)^{3}\kappa\mathsf{cost}({\mathcal{C}}_{\hat{{\mathcal{Y}}}},{\mathcal{S}})+16C_{M}(1+\gamma)^{3}\mathsf{cost}({\mathcal{C}}^{opt}_{\mathcal{S}},{\mathcal{S}})$
$+C_{M}(1+\gamma)^{3}\cdot V^{\prime}(d,k,\varepsilon,T,\Lambda))+(1+\gamma)^{2}V^{\prime\prime}(d,k,\varepsilon,T,\Lambda)$ (29)
We derive the desired expression in the theorem statement by starting with the LHS of Section D.2 and then applying eq. 28 to obtain eq. 32. By observing that the cost of $\hat{{\mathcal{Y}}}_{off}$ wrt its optimal set of centers ${\mathcal{C}}^{opt}_{\hat{{\mathcal{Y}}}}$ is less than its cost wrt to ${\mathcal{C}}^{opt}_{{\mathcal{S}}}$, we obtain eq. 33. eq. 34 is obtained by simply applying the RHS of the coreset relation from Section D.2 wrt ${\mathcal{C}}^{opt}_{{\mathcal{S}}}$. The final statement is derived by grouping like terms and simplifying eq. 35.
Appendix E Missing Proofs from Appendix B
Proof of Proposition 2.